# Weighted Birkhoff Averages Accelerate Data-Driven Methods

### v1

@@ 3.3 Example Method III: Model identification
producing an ODE model of the form $x^{\prime}=\mathbb{X}^{\prime}{\Psi^{\top}}^{\dagger}\vec{\psi}(x)^{\top}$. The sparse identification of nonlinear dynamics (SINDy) method [brunton2016discovering] adds one more idea: sparsity. Rather than fitting all possible terms in the dictionary, SINDy searches for a model that uses as few terms as possible. Let $\Xi=[\xi_{jk}]_{j=1,k=1}^{d,L}$. Given a threshold $\eta>0$, small coefficients are pruned and we solve
@@
This procedure is then iterated: update $\xi\leftarrow\tilde{\Xi}$, re-threshold, and repeat until the coefficients no longer change. Convergence results for SINDy are given in [zhang2019convergence].
@@
Input: Sequential snapshots $\{(X_{n},X_{n}^{\prime})\}_{n=1}^{N}$ of system state and derivative, bump function $w\in\mathcal{W}$, dictionary $\vec{\psi}=(\psi_{1}\,\cdots\,\psi_{L})$, and threshold parameter $\eta>0$.
@@
    Compute the matrices $\mathbb{X}^{\prime}=[X_{1}^{\prime}\ \dots\ X_{N}^{\prime}]$, $\Psi$ defined in (3.6), and $W=\mathrm{diag}(w(0),w(1/N),\dots,w(1))$.
@@
      Compute the matrix $\Xi=[\xi_{jk}]_{jk}=(\mathbb{X}^{\prime}W^{1/2})(\Psi^{\top}W^{1/2})^{\dagger}$.
@@
      Compute the minimizer $\tilde{\Xi}=[\tilde{\xi}_{jk}]_{jk}$ of $\|\mathbb{X}^{\prime}W^{1/2}-\tilde{\Xi}\Psi^{\top}W^{1/2}\|^{2}_{F}$ subject to $\tilde{\xi}_{m,k}=0$ if $|\xi_{m,k}|<\eta$.
@@
      Repeat (b) by replacing $\tilde{\Xi}$ with $\Xi$ until these matrices are the same.
@@
where $u=u(x,t)$ is complex-valued, $t\geq 0$, $x\in[-5,5]$, and periodic boundary conditions are imposed. As discussed in [kevrekidis2005statics], if we initialise with $u(x,0)=2\mathrm{sech}(2(x-x_{0}))$ for some $x_{0}\neq 0$, the resulting soliton oscillates harmonically with frequency 1. This behaviour follows from the Ehrenfest theorem [hall2013quantum]. The left panel of Figure 4 shows $|u(x,t)|$ for $x_{0}=2$, exhibiting this oscillation.
@@
Figure 5 reports the error in the recovered model. We compare: least-squares (3.10) (labelled EDMD), its weighted variant (wtEDMD), SINDy (3.11), and weighted SINDy (wtSINDy). For each method, we measure the norm of the difference between the identified coefficient vector $\Xi$ and the exact coefficients $\Xi_{\mathrm{exact}}=[0\,{-}1\,0\,0\,0\,0]$. The left panel uses a sparsity threshold $\eta=10^{-2}$; the right panel uses $\eta=10^{-4}$.
@@ 3.4 Example Method IV: Spectral measures
Thus $\phi_{\lambda}$ evolves along trajectories by pure exponential growth, decay, or oscillation. Koopman eigenfunctions reveal coherent structures in the dynamics: their level sets can correspond to invariant manifolds [mezic2015applications], isostables [mauroy2013isostables], and other dynamically relevant sets.
@@
In general, there is no reason for $\mathcal{K}$ to have a complete basis of eigenfunctions. Nevertheless, the spectral theorem [conway2019course, Thm. X.4.11] gives a diagonalization of $\mathcal{K}$ using a projection-valued measure $\mathcal{E}$ supported on $\sigma(\mathcal{K})\subset\mathbb{S}^{1}$. This assigns to each Borel subset of $\mathbb{S}^{1}$ an orthogonal projector and yields
@@
The measure $\mathcal{E}$ decomposes the state space into invariant subspaces indexed by subsets of the spectrum, even when actual eigenfunctions are scarce. For a review of the data-driven approximation of Koopman spectral measures, see [colbrook2025introductory, Sec. 6].
Now fix a unit-norm observable $\phi\in L^{2}(\mu)$, $\|\phi\|_{L^{2}(\mu)}=1$. The associated spectral measure of $\mathcal{K}$ with respect to $\phi$ is the probability measure $\nu_{\phi}(A)=\langle\mathcal{E}(A)\phi,\phi\rangle$, supported on $\mathbb{S}^{1}$. Spectral measures carry information not only about Koopman eigenfunctions but also about generalsed eigenfunctions [colbrook2024rigorous, colbrook2025rigged], which in turn illuminate features such as global stability of equilibria [mauroy2016global] and ergodic partitions [budivsic2012applied]. These objects play a central role in nonlinear dynamics [mezic2021koopman].
Since $\sigma(\mathcal{K})\subset\mathbb{S}^{1}$, write $\lambda=\mathrm{e}^{\mathrm{i}\theta}$ and transfer $\nu_{\phi}$ to a measure $\xi_{\phi}$ on the periodic interval $\pi,\pi]$ so that $\xi_{\phi}(B)=\nu_{\phi}(\mathrm{e}^{\mathrm{i}B})$ for measurable subsets $B\subseteq[-\pi,\pi]$. A direct computation using the projection-valued measure gives its Fourier coefficients [colbrook2025rigged]:
@@
The filter $\varphi$ should be close to 1 when $x$ is close to 0 and tapers to 0 near $x=\pm 1$; see [colbrook2024rigorous] for details. The full weighted procedure is summarised in Algorithm 4.
@@
We now apply this method to two-dimensional lid-driven cavity flow at various Reynolds numbers. This flow describes an incompressible viscous fluid in a rectangular cavity with a moving lid. As the Reynolds number increases, the flow undergoes a cascade of bifurcations, exhibiting periodic, quasiperiodic, skew-periodic, [^1] Here, “skew-periodic” denotes a flow with a mixture of (quasi)periodic and chaotic components. and eventually chaotic/turbulent dynamics [shen1991hopf, arbabi2017study]. We use the dataset from [arbabi2017study], which provides $20,000$ snapshots of the vorticity field, sampled at time step $0.1$, for Reynolds numbers $13,000$ (periodic), $16,000$ (quasiperiodic), and $19,000$ (skew-periodic). As our observable $\phi$, we take the total kinetic energy of the flow.
@@
Algorithm 5 Weighted measure-preserving EDMD
@@
    Compute $G=\Psi^{*}W\Psi$ and $A=\Phi^{*}W\Phi$ using the matrices $\Psi$ and $\Phi$ defined in (3.6) with $\vec{\phi}=\vec{\psi}$ and $W=\mathrm{diag}(w(0),w(1/N),\dots,w(1))$.
@@
    Compute the SVD of $G^{-1/2}A^{*}G^{-1/2}=U_{1}\Sigma U_{2}^{*}$.
@@
    Compute $K^{(N)}_{w}=G^{-1/2}U_{2}U_{1}^{*}G^{1/2}$ and $V=G^{-1/2}\hat{V}$.
@@
We can also use these spectral approximations in rigged DMD [colbrook2025rigged] to identify generalised Koopman eigenfunctions. A key ingredient is measure-preserving EDMD [colbrook2023mpedmd], which can likewise incorporate weights; see Algorithm 5. We apply this to the cavity flow data, using only $N=1100$ snapshots. We extract approximate generalised eigenfunctions at angles $\theta$ corresponding to multiples of the base frequency in the periodic flow at $\mathrm{Re}=13,000$, denoted $\theta_{\mathrm{per}}\approx 0.1$. Figure 8 shows the resulting modes for $\theta\in\{\theta_{\mathrm{per}},2\theta_{\mathrm{per}},3\theta_{\mathrm{per}},4\theta_{\mathrm{per}}\}$. Here, weighted and unweighted methods give broadly similar structures.
@@ 3.5 Example Method V: Diffusion forecasts
Our final demonstration of weighted averages concerns nonparametric forecasting. Diffusion forecasting methods [giannakis2019data, berry2015nonparametric, thiede2019galerkin, giannakis2020extraction, berry2015nonparametric2] build predictive models directly from data without first fitting a parametric surrogate model. These methods all rely on diffusion maps [coifman2005geometric, coifman2006diffusion] to construct an effective basis of observables from sampled data.
Following [berry2015nonparametric], we assume that the equally spaced samples $\{X_{n}\}_{n=1}^{N+1}$ come from an underlying stochastic differential equation
@@
We assume that the data $\{X_{n}\}_{n=1}^{N+1}$ are sampled from an invariant (equilibrium) density $p_{\mathrm{eq}}(x)$ of (3.22), i.e., $p_{\mathrm{eq}}(x)=\mathrm{e}^{t\mathcal{L}^{*}}p_{\mathrm{eq}}(x)$ for all $t>0$. Using these data, one applies a variable-bandwidth diffusion maps construction [berry2016variable] to learn eigenfunctions $\phi_{j}$ that are orthonormal in $L^{2}(p_{\mathrm{eq}}^{-1})$. The diffusion forecast [berry2015nonparametric] for the density one sampling interval $\tau$ in the future is then
@@
We follow the setup of [berry2015nonparametric]. We construct a 6-lag time-delay embedding of the Niño-3.4 index and compute the diffusion map eigenfunctions. We train the model on monthly data from January 1920 through December 1999, and we evaluate forecasts on January 2000–December 2013. We use $M=14$ eigenfunctions in (3.23).

### v2

@@ 3.3 Example Method III: Model identification
producing an ODE model of the form $x^{\prime}=\mathbb{X}^{\prime}{\Psi^{\top}}^{\dagger}\vec{\psi}(x)^{\top}$. The sparse identification of nonlinear dynamics (SINDy) method [10] adds one more idea: sparsity. Rather than fitting all possible terms in the dictionary, SINDy searches for a model that uses as few terms as possible. Let $\Xi=[\xi_{jk}]_{j=1,k=1}^{d,L}$. Given a threshold $\eta>0$, small coefficients are pruned and we solve
@@
This procedure is then iterated: update $\xi\leftarrow\tilde{\Xi}$, re-threshold, and repeat until the coefficients no longer change. Convergence results for SINDy are given in [52].
@@
Input: Sequential snapshots $\{(X_{n},X_{n}^{\prime})\}_{n=1}^{N}$ of system state and derivative, bump function $w\in\mathcal{W}$, dictionary $\vec{\psi}=(\psi_{1}\,\cdots\,\psi_{L})$, threshold parameter $\eta>0$, maximum number of iterates $k_{\mathrm{max}}$.
@@
    Compute the matrices $\mathbb{X}^{\prime}=[X_{1}^{\prime}\ \dots\ X_{N}^{\prime}]$, $\Psi$ defined in (3.6), and $W=\mathrm{diag}(w(0),w(1/N),\dots,w((N-1)/N))$.
@@
      Compute set of indices $\mathcal{I}=\{(j,k):|\Xi_{jk}|\geq\eta\}$.
@@
      Set $\Xi_{\mathrm{new}}$ to be the $d\times L$ zero matrix.
@@
      For $j=1,\ldots,d$, set $\mathcal{I}_{j}=\{k:(j,k)\in\mathcal{I}\}$ and update
      $\Xi_{\mathrm{new}}(j,\mathcal{I}_{j})\leftarrow[\mathbb{X}^{\prime}W^{1/2}](j,:)([\Psi^{\top}W^{1/2}](\mathcal{I}_{j},:))^{\dagger}.$
    4. (d)
      If $\Xi_{\mathrm{new}}=\Xi$ or $k\geq k_{\mathrm{max}}$ terminate, otherwise update $\Xi\leftarrow\Xi_{\mathrm{new}}$ and $k\leftarrow k+1$, and return to step (a).
@@
where $u=u(x,t)$ is complex-valued, $t\geq 0$, $x\in[-5,5]$, and periodic boundary conditions are imposed. As discussed in [32], if we initialise with $u(x,0)=2\mathrm{sech}(2(x-x_{0}))$ for some $x_{0}\neq 0$, the resulting soliton oscillates harmonically with frequency 1. This behaviour follows from the Ehrenfest theorem [29]. The left panel of Figure 4 shows $|u(x,t)|$ for $x_{0}=2$, exhibiting this oscillation.
@@
Figure 5 reports the error in the recovered model. We compare: least-squares (3.10) (labelled EDMD), its weighted variant (wtEDMD), SINDy (3.11), and weighted SINDy (wtSINDy). For each method, we measure the Frobenius-norm error $\|\Xi^{(N)}-\Xi_{\mathrm{exact}}\|_{F}$ between the coefficient vector $\Xi^{(N)}$ identified from $N$ snapshots and the exact coefficients $\Xi_{\mathrm{exact}}=[0\,{-}1\,0\,0\,0\,0]$. The left panel uses a sparsity threshold $\eta=10^{-2}$; the right panel uses $\eta=10^{-4}$.
@@ 3.4 Example Method IV: Spectral measures
Thus $\phi_{\lambda}$ evolves along trajectories by pure exponential growth, decay, or oscillation. Koopman eigenfunctions reveal coherent structures in the dynamics: their level sets can correspond to invariant manifolds [38], isostables [35], and other dynamically relevant sets.
@@
In general, there is no reason for $\mathcal{K}$ to have a complete basis of eigenfunctions. Nevertheless, the spectral theorem [22, Thm. X.4.11] gives a diagonalization of $\mathcal{K}$ using a projection-valued measure $\mathcal{E}$ supported on $\sigma(\mathcal{K})\subset\mathbb{S}^{1}$. This assigns to each Borel subset of $\mathbb{S}^{1}$ an orthogonal projector and yields
@@
The measure $\mathcal{E}$ decomposes the state space into invariant subspaces indexed by subsets of the spectrum, even when actual eigenfunctions are scarce. For a review of the data-driven approximation of Koopman spectral measures, see [15, Sec. 6].
Now fix a unit-norm observable $\phi\in L^{2}(\mu)$, $\|\phi\|_{L^{2}(\mu)}=1$. The associated spectral measure of $\mathcal{K}$ with respect to $\phi$ is the probability measure $\nu_{\phi}(A)=\langle\mathcal{E}(A)\phi,\phi\rangle$, supported on $\mathbb{S}^{1}$. Spectral measures carry information not only about Koopman eigenfunctions but also about generalsed eigenfunctions [17, 16], which in turn illuminate features such as global stability of equilibria [36] and ergodic partitions [12]. These objects play a central role in nonlinear dynamics [39].
Since $\sigma(\mathcal{K})\subset\mathbb{S}^{1}$, write $\lambda=\mathrm{e}^{\mathrm{i}\theta}$ and transfer $\nu_{\phi}$ to a measure $\xi_{\phi}$ on the periodic interval $\pi,\pi]$ so that $\xi_{\phi}(B)=\nu_{\phi}(\mathrm{e}^{\mathrm{i}B})$ for measurable subsets $B\subseteq[-\pi,\pi]$. A direct computation using the projection-valued measure gives its Fourier coefficients [16]:
@@
The filter $\varphi$ should be close to 1 when $x$ is close to 0 and tapers to 0 near $x=\pm 1$; see [17] for details. The full weighted procedure is summarised in Algorithm 4.
@@
We now apply this method to two-dimensional lid-driven cavity flow at various Reynolds numbers. This flow describes an incompressible viscous fluid in a rectangular cavity with a moving lid. As the Reynolds number increases, the flow undergoes a cascade of bifurcations, exhibiting periodic, quasiperiodic, skew-periodic, [^1] Here, “skew-periodic” denotes a flow with a mixture of (quasi)periodic and chaotic components. and eventually chaotic/turbulent dynamics [43, 1]. We use the dataset from [1], which provides $20,000$ snapshots of the vorticity field, sampled at time step $0.1$, for Reynolds numbers $13,000$ (periodic), $16,000$ (quasiperiodic), and $19,000$ (skew-periodic). As our observable $\phi$, we take the total kinetic energy of the flow.
@@
Algorithm 5 Weighted measure-preserving EDMD. The use of QR decompositions differs from [20] and improves numerical robustness [19].
@@
    Compute the matrices $\Psi$ and $\Phi$ defined in (3.6) with $\vec{\phi}=\vec{\psi}$ and set $W=\mathrm{diag}(w(0),w(1/N),\dots,w((N-1)/N))$.
@@
    Compute a column-pivoted economy QR decomposition $W^{1/2}\Psi=QRP^{\top}$.
    ($Q$ has orthonormal columns, $R$ is upper triangular with positive diagonals, and $P$ is a permutation matrix)
@@
    Compute the eigendecomposition of the unitary matrix $U_{2}U_{1}^{*}=\hat{V}\Lambda\hat{V}^{*}$.
    (via a Schur decomposition)
  5. 5.
    Compute $K^{(N)}_{w}=PR^{-1}U_{2}U_{1}^{*}RP^{\top}$ and $V=PR^{-1}\hat{V}$.
@@
We can also use these spectral approximations in rigged DMD [16] to identify generalised Koopman eigenfunctions. A key ingredient is measure-preserving EDMD [20], which can likewise incorporate weights; see Algorithm 5. We apply this to the cavity flow data, using only $N=1100$ snapshots. We extract approximate generalised eigenfunctions at angles $\theta$ corresponding to multiples of the base frequency in the periodic flow at $\mathrm{Re}=13,000$, denoted $\theta_{\mathrm{per}}\approx 0.1$. Figure 8 shows the resulting modes for $\theta\in\{\theta_{\mathrm{per}},2\theta_{\mathrm{per}},3\theta_{\mathrm{per}},4\theta_{\mathrm{per}}\}$. Here, weighted and unweighted methods give broadly similar structures.
@@ 3.5 Example Method V: Diffusion forecasts
Our final demonstration of weighted averages concerns nonparametric forecasting. Diffusion forecasting methods [28, 2, 45, 27, 3] build predictive models directly from data without first fitting a parametric surrogate model. These methods all rely on diffusion maps [13, 14] to construct an effective basis of observables from sampled data.
Following [2], we assume that the equally spaced samples $\{X_{n}\}_{n=1}^{N+1}$ come from an underlying stochastic differential equation
@@
We assume that the data $\{X_{n}\}_{n=1}^{N+1}$ are sampled from an invariant (equilibrium) density $p_{\mathrm{eq}}(x)$ of (3.22), i.e., $p_{\mathrm{eq}}(x)=\mathrm{e}^{t\mathcal{L}^{*}}p_{\mathrm{eq}}(x)$ for all $t>0$. Using these data, one applies a variable-bandwidth diffusion maps construction [4] to learn eigenfunctions $\phi_{j}$ that are orthonormal in $L^{2}(p_{\mathrm{eq}}^{-1})$. The diffusion forecast [2] for the density one sampling interval $\tau$ in the future is then
@@
We follow the setup of [2]. We construct a 6-lag time-delay embedding of the Niño-3.4 index and compute the diffusion map eigenfunctions. We train the model on monthly data from January 1920 through December 1999, and we evaluate forecasts on January 2000–December 2013. We use $M=14$ eigenfunctions in (3.23).